# The Curious Case of Factual (Mis)Alignment between LLMs’ Short- and Long-Form Answers

### v2

@@ Abstract
Large language models (LLMs) can correctly answer "When was Einstein born?" yet fail to provide the same date when writing about Einstein’s life—revealing a fundamental inconsistency in how models access factual knowledge across task complexities. While models display impressive accuracy on factual question-answering benchmarks, the reliability gap between simple and complex queries remains poorly understood, eroding their trustworthiness. In this work, we introduce Short-Long Form Alignment for Factual Question Answering (SLAQ), a controlled evaluation framework that compares LLMs’ answers to the same factual questions asked (a) in isolation (short) vs. (b) integrated into complex queries (long). Looking at 16 LLMs across 600 queries, we find a systematic misalignment of answers to the corresponding short and long queries. We further uncover position-dependent accuracy loss and momentum effects where consecutive correct or incorrect answers create self-reinforcing patterns. Through mechanistic analysis, we find that aligned facts activate overlapping model internals, and that metrics based on mechanistic similarity can predict short-long answer alignment with up to 78% accuracy. Our work establishes factual consistency over query complexity as an important aspect of LLMs’ trustworthiness and challenges current evaluation practices, which implicitly assume that good performance for simple factual queries implies reliability in more complex knowledge-seeking tasks too.
@@ 1 Introduction
Studying 16 LLMs through the lens of SLAQ, we find that, while models exhibit substantial short-long alignment w.r.t. factual answer correctness, most of this alignment stems from incorrect answers, i.e., LLMs produce incorrect answers for the same factual question in both long and short-form queries (but it is not necessarily the same answer). We observe that models consistently demonstrate higher factual accuracy in responses to short queries than in long-query responses: the majority of misalignment cases thus stem from a (1) correct answer to the short query and an (2) incorrect answer to the corresponding question included in the long query. Beyond evaluating factual question answering accuracy for both query formats, we identify two critical patterns in model behavior for long-form responses: (1) position-dependent degradation, where factual accuracy declines substantially from 51% for facts appearing early in responses to 30% for facts appearing later, and (2) momentum effects, where consecutive correct answers increase the likelihood of subsequent accuracy, while errors tend to cascade and compound.
To understand the mechanistic basis of the observed factual misalignment, we next analyze the model internals— attention and MLP activation patterns—and identify minimal sets of model components responsible for answer generation for short and long-form queries, respectively. Using zero-ablation Olsson et al. (2022) activation patching Meng et al. (2022), we find that aligned answers activate significantly more similar computational pathways and exhibit stronger correlations in component importance rankings. Moreover, we show that these circuit-level differences have predictive power: employing six pathway similarity metrics, we can predict with 78% accuracy (ROC-AUC: 0.85) whether the answers to the same factual question will align between the two query formats, short and long; here we identify attention head rank correlation as the most predictive feature.
@@ 3.2 Dataset
Figure 3: Long-form QA dynamics by sub-fact position. (a) Slot accuracy: percent correct for each fact position (slots 1–5) in the LQ answer. (b) Trailing 1-streak: $P(\mathrm{correct})$ for the current slot (2–5), conditioned on the length of the immediately preceding run of correct slots. (c) Trailing 0-streak: $P(\mathrm{correct})$ for the current slot (2–5), conditioned on the length of the immediately preceding run of incorrect slots.
@@ 4.1 Experimental Setup
We evaluate models from five families, spanning 1B to 12B parameters: Qwen-2.5 (Yang et al., 2024), Qwen-3 and Qwen-3-Reasoning (Yang et al., 2025a), Llama-3 (Grattafiori et al., 2024), Gemma-2 (Team et al., 2024), and Gemma-3 (Team et al., 2025). All models use greedy decoding via the Hugging Face API. We constrain short-form responses to single sentences and instruct models to provide only requested information for long-form queries (see Table 4 in the Appendix for prompts).
@@ 4.2 Results and Analysis
Position-dependent degradation. In Figure 3a, we observe a monotonic decline in factual accuracy based on the order in which facts are requested in LQ prompts. Accuracy drops from 51.3% for the first requested fact to 30.1% for the fifth, a 21.2 percentage point degradation. Unlike the U-shaped patterns found in input processing (Liu et al., 2024), this query-position effect is strictly linear, with approximately 5% accuracy loss per position. While Yang et al. (2025b) documented hallucinations concentrating at response ends in the summarization task, we find that in open-domain long-form QA, degradation correlates with request order in the prompt itself, suggesting that managing multiple factual requirements imposes a cumulative load that progressively impairs retrieval accuracy.
Momentum within a response. According to Figure 3b and 3c, following consecutive correct answers (positive momentum), accuracy increases from a  30% baseline to 57% after four correct facts: each additional success adds roughly 7% to subsequent accuracy. Conversely, consecutive errors (negative momentum) reduce accuracy from  45% to 24% after three mistakes. This not only confirms Zhang et al. (2024)’s finding of “snowballing” for long-form QA but extends it by quantifying the error propagation and success reinforcement effects. These momentum dynamics, combined with position effects, explain why LQ responses systematically underperform short-form ones even when eliciting the very same factual knowledge.
@@ 5 Mechanistic Analysis
Figure 4: Circuit similarity comparison between aligned and misaligned facts across six metrics. Aligned facts (green) show significantly (p < 0.001) higher mechanistic similarity than misaligned facts (red) for all measures.
@@ 5.1 Preliminaries: Component Importance via Zero-Ablation
Figure 5: Predictive modeling performance using circuit similarity metrics. (a) Individual feature performance shows Spearman Attention as the strongest single predictor of factual alignment (ROC-AUC = 0.83). (b) Combined features achieve robust performance across all evaluation metrics (ROC-AUC = 0.81, Accuracy = 0.76). (c) Feature importance reveals Spearman Attention as the dominant predictor (coefficient = 1.36).
@@ 5.2 Mechanistic Comparison
Experimental Setup We analyze four models spanning different scales: Qwen-2.5 (1B, 3B) and Qwen-3 (1.7B, 4B). For each model, we select 60 short-long fact pairs that are divided into 30 correctly aligned pairs and 30 misaligned pairs. This yields us 240 short-long fact pairs. For each fact, we obtain: (1) Minimal component sets for all answer tokens (2) Importance scores for all attention heads and MLP layers. We set the the greedy threshold to 90% are are able to recover the original token with 100% accuracy.
Results. Figure 4 reveals that aligned facts exhibit systematically higher mechanistic similarity than misaligned facts across all six metrics (p < 0.05 for all comparisons). This provides direct evidence that behavioral alignment reflects distinct internal mechanisms. Node IoU shows the largest relative difference, with aligned facts achieving 16.2% higher similarity (0.316 vs 0.272). Containment shows a 10.6% increase (0.552 vs 0.499). These gaps suggest that consistent responses recruit overlapping computational pathways, while inconsistent responses activate different mechanisms. Spearman correlations are consistently higher than Pearson correlations for both attention (0.909 vs 0.744) and MLP components (0.778 vs 0.673), suggesting that components’ rank by importance is more relevant than their raw activation magnitudes. In other words, factual consistency depends on maintaining stable computational pathways, not identical activation patterns.
@@ 5.3 Model Internals as Predictors
Building on our finding that aligned facts exhibit significantly higher mechanistic similarity, we investigate whether these similarity metrics can predict factual alignment. We train a logistic regression classifier using the six similarity metrics as features on a dataset of 240 fact pairs (120 aligned, 120 misaligned) across four models (Qwen-2.5 1B/3B, Qwen-3 1.7B/4B). We evaluate performance via 5-fold cross-validation to assess both individual feature contributions and combined predictive power.
Results and Analysis. The predictive modeling results in Figure 5 validate the mechanistic basis of factual alignment and reveal which similarity metrics best capture this phenomenon. Spearman correlation over attention components emerges as the strongest individual predictor of alignment (ROC-AUC = 0.83), indicating that attention importance hierarchies provide the most reliable measurable signal of factual consistency. The combined feature model achieves robust performance (ROC-AUC = 0.81, Accuracy = 0.76), with logistic regression coefficients revealing the underlying computational logic: Spearman attention dominates with a coefficient of 1.36, whereas Pearson correlation gets a negative coefficient: this suggests that linear magnitude similarities actually predict misalignment. This convergence in findings between (i) similarity analysis and (ii) predictive modeling shows that mechanistic interpretability can offer not just explanatory insights into model behavior but also practical tools for predicting inconsistencies in LLMs responses for the same factual questions.
@@ 6 Conclusion
This work establishes factual consistency over query complexity as a critical dimension of LLM reliability. Through SLAQ, we demonstrate that LLMs exhibit systematic misalignment when answering identical factual questions embedded in queries of varying complexity. Our analysis reveals predictable patterns: factual accuracy degrades substantially with response position, declining from 51% for early facts to 30% for later ones, and correctness exhibits momentum effects. Our mechanistic analysis provides the first empirical evidence that behavioral factual alignment corresponds to similar internal mechanisms. Through zero-ablation, we found that aligned facts exhibit higher mechanistic similarity. The predictive modeling results demonstrate practical applications, with mechanistic similarity metrics achieving 78% accuracy in predicting factual alignment. Future work should investigate targeted interventions on circuit patterns to address the consistency failures identified in LLMs. We encourage future research to benchmark their models on SLAQ dataset to test for factual alignment.
@@
SLAQ Dataset and Framework The SLAQ dataset and framework have several limitations. (1) The dataset is synthetically generated. While a fully human-annotated dataset would be ideal, it is cost-prohibitive; using LLMs as data generators offers a more balanced quality–resource trade-off. We mitigate this limitation by manually verifying each prompt–answer pair and find that OpenAI’s o3-mini-high is a strong data generator for the SLAQ task (§3.2). (2) The dataset—and this work—currently covers only English. (3) For evaluation, we adopt the LLM-as-a-judge paradigm to assess factual correctness; in our experiments, however, we find Gemini-2.5-Flash achieves high agreement with human annotations (see §4).

### v3

@@ Abstract
Large language models (LLMs) can correctly answer "When was Einstein born?" yet fail to provide the same date when the same question is embedded in a long-form query requesting multiple facts about Einstein’s life, revealing a fundamental inconsistency in how models access facts across different complexities of knowledge seeking tasks. While models display impressive accuracy on factual question-answering benchmarks, the reliability gap between simple and complex long-form queries remains poorly understood, eroding their trustworthiness. In this work, we introduce Short-Long Form Alignment for Factual Question Answering (SLAQ), a controlled evaluation framework that compares LLMs’ answers to the same factual questions asked (a) in isolation (short) vs. (b) integrated into complex queries (long). Looking at 16 LLMs across 600 queries, we find a systematic misalignment of answers to the corresponding short and long queries. We further uncover momentum effects where consecutive correct or incorrect answers create self-reinforcing patterns. Through mechanistic analysis, we find that aligned facts activate overlapping model internals, and that metrics based on mechanistic similarity can predict short-long answer alignment with up to 80% accuracy. Our work establishes factual consistency over query complexity as an important aspect of LLMs’ trustworthiness and challenges current evaluation practices, which implicitly assume that good performance for simple factual queries implies reliability in more complex knowledge-seeking tasks as well.
@@ 1 Introduction
Studying 16 LLMs through the lens of SLAQ, we find that, while models exhibit substantial short-long alignment w.r.t. factual answer correctness, most of this alignment stems from incorrect answers, i.e., LLMs produce incorrect answers for the same factual question in both long and short-form queries (but it is not necessarily the same answer). We observe that models consistently demonstrate higher factual accuracy in responses to short queries than in long-query responses: the majority of misalignment cases thus stem from a (1) correct answer to the short query and an (2) incorrect answer to the corresponding question included in the long query. Beyond evaluating factual question answering accuracy for both query formats, we identify momentum effects, where consecutive correct answers increase the likelihood of subsequent accuracy, while errors tend to cascade and compound.
To understand the mechanistic basis of the observed factual misalignment, we next analyze the model internals— attention and MLP activation patterns—and identify minimal sets of model components responsible for answer generation for short and long-form queries, respectively. Using zero-ablation Olsson et al. (2022) activation patching Meng et al. (2022), we find that aligned answers activate significantly more similar computational pathways and exhibit stronger correlations in component importance rankings. Moreover, we show that these circuit-level differences have predictive power: employing six pathway similarity metrics, we can predict with 80% accuracy (ROC-AUC: 0.85) whether the answers to the same factual question will align between the two query formats, short and long; here we identify attention head rank correlation as the most predictive feature.
@@ 3.2 Dataset
Figure 3: Long-form QA dynamics by sub-fact position. (a) Trailing 1-streak: $P(\mathrm{correct})$ for the current slot (2–5), conditioned on the length of the immediately preceding run of correct slots. (b) Trailing 0-streak: $P(\mathrm{correct})$ for the current slot (2–5), conditioned on the length of the immediately preceding run of incorrect slots. Slots (1-5) are sub-fact positions in the long-form query.
@@ 4.1 Experimental Setup
We evaluate models from five families, spanning 1B to 12B parameters: Qwen-2.5 (Yang et al., 2024), Qwen-3 and Qwen-3-Reasoning (Yang et al., 2025a), Llama-3 (Grattafiori et al., 2024), Gemma-2 (Team et al., 2024), and Gemma-3 (Team et al., 2025). All models use greedy decoding via the Hugging Face API. We constrain short-form responses to single sentences and instruct models to provide only requested information for long-form queries (see Table 4 in the Appendix for prompts). To control for positional effects, we randomize the order of the five sub-questions within each long-form query across 5 permutations and report averaged results.
@@ 4.2 Results and Analysis
Momentum within a response. According to Figure 3b and 3c, following consecutive correct answers (positive momentum), accuracy increases from a  30% baseline to 57% after four correct facts: each additional success adds roughly 7% to subsequent accuracy. Conversely, consecutive errors (negative momentum) reduce accuracy from  45% to 24% after three mistakes. This not only confirms Zhang et al. (2024)’s finding of “snowballing” for long-form QA but extends it by quantifying the error propagation and success reinforcement effects. These momentum dynamics explain why LQ responses systematically underperform short-form ones even when eliciting the very same factual knowledge.
@@ 5 Mechanistic Analysis
Figure 4: Mechanistic similarity comparison between aligned and misaligned facts across six metrics averaged across six LLMs. Aligned facts (green) show significantly (p < 0.05) higher mechanistic similarity than misaligned facts (red) for all measures. Scores are averaged across six LLMs.
@@ 5.1 Preliminaries: Component Importance via Zero-Ablation
Figure 5: Predictive modeling performance using circuit similarity metrics. (a) Individual feature performance shows Spearman Attention as the strongest single predictor of factual alignment (ROC-AUC = 0.86). (b) Combined features achieve robust performance across all evaluation metrics (ROC-AUC = 0.85, Accuracy = 0.80). (c) Feature importance reveals Spearman Attention as the dominant predictor (coefficient = 1.53).
@@ 5.2 Mechanistic Comparison
Experimental Setup We analyze four models spanning different scales: Qwen-2.5 (1B, 3B) and Qwen-3 (1.7B, 4B). For each model, we select 60 short-long fact pairs that are divided into 30 correctly aligned pairs and 30 misaligned pairs. This yields us 240 short-long fact pairs. For each fact, we obtain: (1) Minimal component sets for all answer tokens (2) Importance scores for all attention heads and MLP layers. We set the the greedy threshold to 90% are are able to recover the original token with 100% accuracy. After extracting minimal components and importance scores, we compute pairwise similarity measures between all tokens in the short answer and the corresponding-fact tokens in the long answer span, then apply EMD to aggregate these into a single answer-level similarity score.
Results. Figure 4 reveals that aligned facts exhibit significantly higher mechanistic similarity than misaligned facts across all six metrics (all p < 0.001). This yields direct evidence that behavioral alignment reflects distinct internal mechanisms.
Set-based metrics show substantial differences: IoU increases by 18.4% for aligned facts (0.309 vs 0.261), while Containment shows a 12.0% increase (0.552 vs 0.493). These gaps indicate that factually consistent responses recruit more overlapping components, while inconsistent responses activate divergent computational pathways. Correlation-based metrics reveal complementary patterns. Pearson correlations achieve higher absolute values than Spearman correlations for both attention (0.748 vs 0.686) and MLP components (0.918 vs 0.796), indicating strong linear relationships in raw activation magnitudes. Examining discriminative power, attention metrics show the largest relative gaps regardless of correlation type (Spearman: 14.1%, Pearson: 14.2%), suggesting attention patterns are most diagnostic of factual alignment. In contrast, MLP components show divergent patterns: Spearman correlations provide stronger discrimination (8.9%) than Pearson correlations (3.3%), indicating that the ranking of MLPs by importance is more informative than their raw activation (i.e., importance) magnitudes.
Overall, these results confirm our hypothesis: factual alignment corresponds to mechanistic similarity, with aligned facts recruiting more similar computational pathways.
@@ 5.3 Model Internals as Predictors
Building on our finding that aligned facts exhibit significantly higher mechanistic similarity, we investigate whether these similarity metrics can predict factual alignment. We train a logistic regression classifier using the six similarity metrics as features on a dataset of 360 fact pairs (180 aligned, 180 misaligned) across six models (Llama-3.2 1B/3B, Qwen-2.5 1B/3B, Qwen-3 1.7B/4B). We evaluate performance via 5-fold cross-validation to assess both individual feature contributions and combined predictive power.
Results and Analysis. The predictive modeling results in Figure 5 validate the mechanistic basis of factual alignment and reveal which similarity metrics best capture this phenomenon. Spearman attention emerges as the strongest individual predictor of alignment (ROC-AUC = 0.86, Accuracy = 0.79), indicating that ranks of attention components by importance provide the most reliable measurable signal of factual consistency. The combined feature model achieves robust performance (ROC-AUC = 0.85, Accuracy = 0.80), with logistic regression coefficients revealing the underlying computational structure: Spearman Attention dominates with a coefficient of 1.53, more than four times larger than the next most discriminative feature (Containment).
These results demonstrate that mechanistic interpretability offers not just explanatory insights into model behavior but also a practical tool for predicting factual inconsistencies in LLMs’ responses..
@@ 6 Conclusion
This work establishes factual consistency over query complexity as a critical dimension of LLM reliability. Through SLAQ, we demonstrate that LLMs exhibit systematic misalignment when answering identical factual questions embedded in queries of varying complexity. Our analysis reveals that LLMs are mostly aligned on being factually incorrect and correctness exhibits momentum effects. Our mechanistic analysis provides the first empirical evidence that behavioral factual alignment corresponds to similar internal mechanisms. The predictive modeling results demonstrate practical applications in detecting factual misalignment. Future work should explore factual alignment between short and free-form long-form queries, and investigate targeted interventions on circuit patterns to address the consistency failures identified in LLMs.
@@
SLAQ Dataset and Framework The SLAQ dataset and framework have several limitations. (1) The dataset is synthetically generated. While a fully human-annotated dataset would be ideal, it is cost-prohibitive; using LLMs as data generators offers a more balanced quality–resource trade-off. We mitigate this limitation by manually verifying each prompt–answer pair and find that OpenAI’s o3-mini-high is a strong data generator for the SLAQ task (§3.2). (2) The dataset—and this work—currently covers only English. (3) For evaluation, we adopt the LLM-as-a-judge paradigm to assess factual correctness; in our experiments, however, we find Gemini-2.5-Flash achieves high agreement with human annotations (see §4). (4) Our dataset and evaluation framework does not address free-form long-form generation.